Joe and Jane together have $91. If Joe has four-ninths as much as Jane has, who has more money and how much more?
Answer by stanbon(75887)

Joe and Jane together have $91. If Joe has four-ninths as much as Jane has, who has more money and how much more?
------
Equations::
o + a = 91
o = (4/9)a
--------
Substitute for "o" and solve for "a":
(4/9)a + a = 91
---------
(13/9)a = 91
----------
a = (9/13)91
----------
a = 9*7
--------
a = 63 yrs (Jane's age)
----------------
Solve for "o"
o = 91-63
o = 28 yrs (Joe's age)
-----------------------
Cheers,
Stan H.
---------------